OpenStudy (anonymous):

2 5 7 12 19 31 50 __________ a) 81 b) 60 c) 70 d) 84

OpenStudy (larseighner):

Try writing the difference between each term and the one that follows it.

OpenStudy (anonymous):

81

OpenStudy (anonymous):

here 7=5+2 12 =7+5 19=12+7 and so on...

OpenStudy (larseighner):

That's right.

OpenStudy (anonymous):

Adding by last 2 numbers will continue the series? right?

OpenStudy (anonymous):

and it applys on all questions like this one?

OpenStudy (larseighner):

Right. You cannot tell where the first two numbers came from, but they give you enough numbers to figure out the following ones.

OpenStudy (anonymous):

1 2 2 4 8 32 __________ a)256 b) 629 c) 430 d) 512

OpenStudy (anonymous):

but in this one multiplying the last 2 figures is the right answer.. my question is how we figure out the difrnc of adding or multiplying?

OpenStudy (anonymous):

yup

OpenStudy (larseighner):

In general you can't. You have to try out one, and if that doesn't work, try out the other.

OpenStudy (anonymous):

and whts the answer of this one? 49 36 25 16 9 4 __________ a) 1 b) 2 c) 3 d) 4

OpenStudy (anonymous):

1

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

all are squares of natural numbers in decreasing order square of 7 is 49 square of 6 is 36 square of 5 is 25 and so on..

OpenStudy (anonymous):

ohhh.. ok i got it..:D